A 1litre vessel which is equipped with a movable piston is filled with a 2M aqueous solution of h2O2.the h2O2 decomposes to h2O and O2 is a first order process with half life 5 hours at 300K.as gas formed the piston moves upward against external pressure of 1atm.calculate work done by the gas from the start of 6hours till the end of 15 hours
1
Expert's answer
2019-05-23T08:00:42-0400

2H2O2 = 2H2O + O2;

t 1/2 = ln2/K;

K = ln2/ t 1/2 = 0.693/5 = 0.14 (1/hours);

K = 1/t * ln(C0/C);

lnC = lnC0/(K*t)

lnC = 0.693/(0.14*9) = 0.55;

C = e^0.87 = 1.7